OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (anonymous):

a1 x+b1y=c1 a2 x+b2y=c2 \[if \frac{ a1 }{a2}=\frac{ b1 }{ b2 }=\frac{ c1 }{ c2 },then infinite solutions.\] \[if \frac{ a1 }{a2 }=\frac{ b1 }{ b2 }\neq \frac{ c1 }{ c2 },then no solution.\]